OpenStudy (anonymous):

Ax^2 + Dx + Ey + F = 0; A doesn't equal 0 is a vertical line if D^2=4AF and E=0 i need points to graph

OpenStudy (anonymous):

If E is zero, this is a quadratic equation. D^2 = 4AF indicates there is only one solution.

OpenStudy (anonymous):

That solution will be x = -D/2A.

OpenStudy (anonymous):

okay so put in the quadratic equation thing

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

Facts to remember about this problem: \[ax^2+bx+c=0 \implies x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Note that the conditions given in the problem make the discriminant (the argument of the square root in the fomula) zero.

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

so do D=2 square root AF

OpenStudy (anonymous):

I suppose x=-b/2a could be considered a vertical line in the Cartesian plane, but it is usually thought of as a point on the real line.

OpenStudy (anonymous):

ohhhhh nevermind i got it

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

so if D^2 > 4AF what happens

OpenStudy (anonymous):

Then there are two distinct, real-number solutions.

OpenStudy (anonymous):

You can get them from the quadratic formula.

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

what is D^2 < 4AF

OpenStudy (anonymous):

@AnimalAin

OpenStudy (anonymous):

In that case, there is no real-number solution. The two solutions are complex-conjugates of the form a +bi and a - bi.